# Lightning-Fast Image Inversion and Editing for Text-to-Image Diffusion Models

### v4

@@ 1 Introduction
As inversion became a critical building block in various tasks, several inversion methods have been suggested. Denoising Diffusion Implicit Models (DDIM) (Song et al., 2021) introduced a deterministic and fast sampling technique for image generation with diffusion models. DDIM inversion transforms an image back into a latent noise representation by approximating the inversion equation. Although this approximation makes it very fast, it also introduces an approximation error (as explained in section 3), causing noticeable distortion artifacts in the reconstructed images. This is particularly noticeable in few-step diffusion models (Luo et al., 2023; Sauer et al., 2023; Esser et al., 2024), with large gaps between the diffusion time-steps, and where inference is achieved with only 2 – 4 steps. Several efforts have been made to address inconsistencies in DDIM Inversion, which result in poor reconstruction quality  (Mokady et al., 2023; Wallace et al., 2022; Pan et al., 2023). For example, (Pan et al., 2023; Garibi et al., 2024) used fixed-point iterations to solve the DDIM-inversion equation. Another approach in (Hong et al., 2024b) tackled the minimization of residual error in DDIM-inversion via gradient descent. Although these methods demonstrate improvements over previous approaches, their editing quality and computational speed are still limited.
@@
In this paper, we frame the deterministic diffusion inversion problem as finding the roots of an implicit function. We propose a solution based on the Newton-Raphson (NR) numerical scheme (Burden et al., 2015) - a very fast and well-tested optimization method for root finding. We further define a scalar NR scheme for this problem that can be computed efficiently. However, NR has a main disadvantage in our context. When applied to our highly non-convex function, it tends to find roots that drift outside the distribution of latents that the diffusion model was trained on. To address this, we introduce a method to guide the NR iterations towards in-distribution roots. This is done by incorporating prior knowledge about the noise distribution added at each diffusion step. This approach enables rapid inversion while maintaining state-of-the-art reconstruction and editing quality.
@@
We name our approach GNRI, for Guided Newton Raphson Inversion. GNRI converges in just a few iteration steps at each diffusion time step and achieves high-quality image inversions (in terms of reconstruction accuracy). In practice, 1-2 iterations are sufficient for convergence that yields significantly more accurate results than other inversion methods. Importantly, GNRI requires no model training, fine-tuning, prompt optimization, or additional parameters, and is compatible with all pre-trained diffusion models and deterministic schedulers. We demonstrate its effectiveness for inverting different deterministic schedulers used with latent diffusion model (Stable Diffusion) (Rombach et al., 2022), few-step latent diffusion model (SDXL-turbo) (Sauer et al., 2023) and few-step flow matching model (Flux.1) (Black-Forest, 2024). Figure 1 demonstrates the quality and speed of GNRI for iterative editing using few-step diffusion models. Using such models models that require 4 denoising steps, our approach can edit real images within 0.4 seconds. This allows users to edit images on the fly, using text-to-image models.
@@ 2 Related work
Text-to-image diffusion models (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; Balaji et al., 2022; Esser et al., 2024) translate random samples (seeds) from a high-dimensional space, guided by a user-supplied text prompt, into corresponding images. DDIM (Song et al., 2021) is a widely used deterministic scheduler, that demonstrates the inversion of an image to its latent noise seed. When applied to the inversion of text-guided diffusion models, DDIM inversion suffers from low reconstruction accuracy that is reflected in further tasks, particularly when applied to few-step diffusion models that require only 3-4 denoising steps. This happens because DDIM inversion relies on a linear approximation, causing the propagation of errors that result in inaccurate image reconstruction. Recent studies (Mokady et al., 2023; Wallace et al., 2022; Pan et al., 2023; Hong et al., 2024b) address this limitation. Null-text inversion (Mokady et al., 2023) optimizes the embedding vector of an empty string. This ensures that the diffusion process calculated using DDIM inversion, aligns with the reverse diffusion process. (Miyake et al., 2023) replace the null-text embedding with a prompt embedding instead. This enhances convergence time and reconstruction quality but results in inferior image editing performance. In both (Mokady et al., 2023) and (Miyake et al., 2023), the optimized embedding must be stored, resulting in nearly 3 million additional parameters for each image (using 50 denoising steps of StableDiffusion  (Rombach et al., 2022)).
@@
EDICT (Wallace et al., 2022) introduced invertible neural network layers, specifically Affine Coupling Layers, to calculate both backward and forward diffusion paths. While effective, it comes at the cost of prolonging inversion time. BDIA (Zhang et al., 2023) introduced a novel approximation tailored for EDICT, enhancing its computational efficiency while maintaining the accuracy in diffusion inversion. However, it still demands considerably more time, approximately ten times longer than DDIM Inversion. TurboEdit (Wu et al., 2024) introduced an encoder-based iterative inversion technique for few-step diffusion models, enabling precise image inversion and disentangled image editing. DirectInv (Ju et al., 2023) separates the source and target diffusion branches for improved text-guided image editing, leading to better content preservation and edit fidelity. AIDI (Pan et al., 2023) and ReNoise (Garibi et al., 2024) used a fixed-point iteration technique at each inversion step to find a better solution for the implicit function posed by DDIM equations and thereby achieving improved accuracy. ExactDPM (Hong et al., 2024b), similar to  (Pan et al., 2023; Garibi et al., 2024) proposed gradient-based methods for finding effective solutions for the implicit function.
@@
Alternative methods proposed in (Huberman-Spiegelglas et al., 2023; Brack et al., 2024; Deutch et al., 2024) use a stochastic DDPM inversion instead of a deterministic one. Although it can recover the exact image during reconstruction, the stochastic nature of these approaches usually causes the method to struggle with reconstructing fine details during editing. These methods require many steps for high-quality editing, making the process time-consuming. Additionally, they demand a larger memory footprint, as they need to store $T+1$ latents for each inverted image, restricting their use only to reconstruction and editing tasks. We also compare our approach with these methods in terms of editing quality and demonstrate that our method outperforms them all.
@@ 3 Preliminaries
$z_{t}$ $=$ $f(z_{t})$ (4)
$f(z_{t})$ ::=:= $\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}z_{t-1}+\sqrt{\alpha_{t}}\cdot\Delta\psi(\alpha_{t})\cdot\epsilon_{\theta}(z_{t},t,p).$
@@
DDIM inversion is fast, but the approximation of Eq. (5) inherently introduces errors at each time step. As these errors accumulate, they cause the whole diffusion process to become inconsistent in the forward and the backward processes, leading to poor image reconstruction and editing (Mokady et al., 2023; Wallace et al., 2022; Pan et al., 2023). This is particularly noticeable in few-step and consistency models with a small number of denoising steps (typically 2-4 steps), where there’s a significant gap between $z_{t}$ and $z_{t-1}$ Garibi et al. (2024).
@@
Iterative inversion and optimization methods. Several papers proposed to improve the approximation using iterative methods. AIDI (Pan et al., 2023) and ReNoise (Garibi et al., 2024) proposed to directly solve Eq. (3) using fixed-point iterations (Burden, 1985), a widely-used method in numerical analysis for solving implicit functions. In a related way, (Hong et al., 2024b) solves a more precise inversion equation, obtained by employing higher-order terms, using gradient descent.
Newton-Raphson Root finding. The Newton-Raphson method is a widely used numerical technique for finding roots of a real-valued function $\mathcal{F}$ (Burden et al., 2015). It is particularly effective for solving equations of the form $\mathcal{F}(z)=0$, when $\mathcal{F}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ for an arbitrary dimension $D$. It provides fast convergence, typically quadratic, by requiring the evaluation of the function and the inversion of its Jacobian matrix. It has also been shown that when initialized near a local extremum, NR may converge to that point (oscillating around it) (Kaw et al., 2003). The NR scheme in general form, is given by
$z^{k+1}_{t}=z^{k}_{t}-J(z^{k}_{t})^{-1}\leavevmode\nobreak\ \mathcal{F}(z^{k}_{t}),$ (6)
where $J(z_{t}^{k})^{-1}$ presents the inverse of a Jacobian matrix $J\in\mathbb{R}^{D\times D}$ for $\mathcal{F}$ (all derivatives are w.r.t to $z$), and $k$ stands for the iteration number. The iteration starts with an initial guess, $z=z^{0}$.
@@ 4 Our method: Guided Newton Raphson Inversion
Existing inversion methods are either fast-but-inaccurate, like DDIM inversion, or more precise-but-slow, like (Pan et al., 2023) and (Hong et al., 2024b). This paper presents a method that achieves a balance of speed and improved precision compared to these existing approaches.
@@ 4.1 Framing inversion as efficient root-finding
and seek the roots $\hat{r}(z_{t})=0$ where |||||$\cdot$||| | ⋅ | | denotes a $L_{1}$ norm which simply sum over all absolute values of $\hat{r}$. By reducing the problem from $r$ to $\hat{r}$, the Jacobian becomes a vector and can be computed quickly (see derivation in Appendix A). We call this method Newton-Raphson Inversion, NRI.
@@ 4.2 Guided Newton Raphson Inversion
where $\mu_{t}$, $\beta_{t}$ were defined in Section 3, see Eq (1). To guide NR to select an ”in-distribution” solution, we define a new objective:
@@
Here, $\lambda>0$ is a hyperparameter weighting factor for the guidance term. See Appendix D for ablation of $\lambda$ values.
@@
Here $g_{i}:=\frac{\partial\mathcal{F}(z_{t})}{\partial z_{t}(i)}$ is the partial derivative of $\mathcal{F}$ with respect to the variable (descriptor) $z_{t}(i)$. The index $i\in[0,1,...,D]$ indicates $z$’s components and $k$ stands for the iteration number, while $\eta$ is a small constant added for numerical stability. The function $g$ can be computed efficiently using automatic differentiation engines. We initialize the process with $z$ from the previous diffusion time-step $z_{t-1}$.
We call our approach GNRI for Guided Newton Raphson Inversion.
@@ 5 Experiments
Compared Methods: We compared our approach with the following methods. Standard DDIM Inversion (Song et al., 2021), SDEdit Meng et al. (2022), Null-text (Mokady et al., 2023) and EDICT (Wallace et al., 2022). Two fixed-point based methods: AIDI (Pan et al., 2023) and ReNoise (Garibi et al., 2024). Gradient-based method ExactDPM (Hong et al., 2024b). Wu2024Turbo (Wu et al., 2024) and two DDPM inversion methods: EditFriendly Huberman-Spiegelglas et al. (2023) and Gal2024Turbo (Deutch et al., 2024).
In all experiments, we used code and parameters provided by the respective authors.
@@ 5.1 Guidance Term Motivation and Comparison to Other Numerical Schemes
To illustrate the motivation behind our guidance term, we compare the reconstruction quality of non-guided NRI a ”vanilla” NR inversion method with two iterative numerical schemes: Fixed-point Iteration (FPI) and Gradient Descent (GD). We test the reconstruction of 5,000 image-caption pairs from COCO Lin et al. (2014), presenting three metrics in Fig. 3.
Fig 3a shows that FPI, GD and non-guided NRI, all converge to low residual values but the reconstruction values for NRI (green line) in Fig. 3b are notably lower. This suggests that solutions reached by NRI are in areas of the latent space that are not processed correctly by the diffusion model.
@@
Figure 4: (Left) Reconstruction qualitative results: Comparing image inversion-reconstruction performance. While all baseline methods struggle to preserve the original image, GNRI successfully excels in accurately reconstructing it. (Right) Inversion Results: Mean reconstruction quality (y-axis, PSNR) and runtime (x-axis, seconds) on the COCO2017 validation set. Our method achieves high PSNR while reducing inversion-reconstruction time by a factor of ×$\times 2$ to ×$\times 40$ on SDXL-turbo and ×$\times 10$ to ×$\times 40$ on Flux.1, compared to other approaches.
@@ 5.2 Image Reconstruction
To evaluate the fidelity of our approach, we measure the PSNR of images reconstructed from seed inversions. Specifically, we used the entire set of 5000 images from the MS-COCO-2017 validation dataset (Lin et al., 2014), along with their corresponding captions. For each image-caption pair, we first found the inverted latent and then used it to reconstruct the image using the same caption. Given that the COCO dataset provides multiple captions for each image, we used the first listed caption as a conditioning prompt. Figure 4 (left) demonstrates a qualitative comparison of reconstructed images. Figure 4 (right) illustrates the PSNR of reconstructed images versus inversion time, highlighting the performance of our method compared to SoTA inversion techniques. The dashed black line represents the upper bound set by the Stable Diffusion VAE, showing that our method achieves the highest PSNR among SoTA methods and approaches the VAE upper bound, while also being the fastest. Additional qualitative comparisons for SD2.1 and further analysis are in Appendix G.

### v5

@@ 1 Introduction
As inversion became a critical building block in various tasks, several inversion methods have been suggested. Denoising Diffusion Implicit Models (DDIM) (Song et al., 2021) introduced a deterministic and fast sampling technique for image generation with diffusion models. DDIM inversion transforms an image back into a latent noise representation by approximating the inversion equation. Although this approximation makes it very fast, it also introduces an approximation error (as explained in section 3), causing noticeable distortion artifacts in the reconstructed images. This is particularly noticeable in few-step diffusion models (Luo et al., 2023; Sauer et al., 2023; Esser et al., 2024), with large gaps between the diffusion time-steps, and where inference is achieved with only 2 – 4 steps. Several efforts have been made to address inconsistencies in DDIM Inversion, which result in poor reconstruction quality  (Mokady et al., 2023; Wallace et al., 2023; Pan et al., 2023). For example, (Pan et al., 2023; Garibi et al., 2024) used fixed-point iterations to solve the DDIM-inversion equation. Another approach in (Hong et al., 2024) tackled the minimization of residual error in DDIM-inversion via gradient descent. Although these methods demonstrate improvements over previous approaches, their editing quality and computational speed are still limited.
@@
In this paper, we frame the deterministic diffusion inversion problem as finding the roots of an implicit function. We propose a solution based on the Newton-Raphson (NR) numerical scheme (Burden et al., 2015) - a very fast and well-tested optimization method for root finding. We further define a scalar NR scheme for this problem that can be computed efficiently. However, NR has a main disadvantage in our context. When applied to our highly non-convex function, it tends to find roots that drift outside the distribution of latents that the diffusion model was trained on.To mitigate this, we introduce a guidance term that leverages prior knowledge of the likely solution locations, steering the NR iterations toward in-distribution roots. This is done by adding this prior knowledge on the noise distribution at each diffusion step. Ultimately, our approach enables rapid inversion while maintaining state-of-the-art reconstruction and editing quality.
@@
We name our approach GNRI, for Guided Newton Raphson Inversion. GNRI converges in just a few iteration steps at each diffusion time step and achieves high-quality image inversions (in terms of reconstruction accuracy). In practice, 1-2 iterations are sufficient for convergence that yields significantly more accurate results than other inversion methods. Importantly, GNRI requires no model training, fine-tuning, prompt optimization, or additional parameters, and is compatible with all pre-trained diffusion and flow-matching models with deterministic schedulers. We demonstrate its effectiveness for inverting different deterministic schedulers used with latent diffusion model (Stable Diffusion) (Rombach et al., 2022), few-step latent diffusion model (SDXL-turbo) (Sauer et al., 2023) and few-step flow matching model (Flux.1) (Black-Forest, 2024). Figure 1 demonstrates the quality and speed of GNRI for iterative editing using few-step diffusion models. Using such models models that require 4 denoising steps, our approach can edit real images within 0.4 seconds. This allows users to edit images on the fly, using text-to-image models.
@@ 2 Related work
Text-to-image diffusion models (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; Balaji et al., 2022; Esser et al., 2024) translate random samples (seeds) from a high-dimensional space, guided by a user-supplied text prompt, into corresponding images. DDIM (Song et al., 2021) is a widely used deterministic scheduler, that demonstrates the inversion of an image to its latent noise seed. When applied to the inversion of text-guided diffusion models, DDIM inversion suffers from low reconstruction accuracy that is reflected in further tasks, particularly when applied to few-step diffusion models that require only 3-4 denoising steps. This happens because DDIM inversion relies on a linear approximation, causing the propagation of errors that result in inaccurate image reconstruction. Recent studies (Mokady et al., 2023; Wallace et al., 2023; Pan et al., 2023; Hong et al., 2024) address this limitation. Null-text inversion (Mokady et al., 2023) optimizes the embedding vector of an empty string. This ensures that the diffusion process calculated using DDIM inversion, aligns with the reverse diffusion process. (Miyake et al., 2023) replace the null-text embedding with a prompt embedding instead. This enhances convergence time and reconstruction quality but results in inferior image editing performance. In both (Mokady et al., 2023) and (Miyake et al., 2023), the optimized embedding must be stored, resulting in nearly 3 million additional parameters for each image (using 50 denoising steps of StableDiffusion  (Rombach et al., 2022)).
@@
EDICT (Wallace et al., 2023) introduced invertible neural network layers, specifically Affine Coupling Layers, to calculate both backward and forward diffusion paths. While effective, it comes at the cost of prolonging inversion time. BDIA (Zhang et al., 2023) introduced a novel approximation tailored for EDICT, enhancing its computational efficiency while maintaining the accuracy in diffusion inversion. However, it still demands considerably more time, approximately ten times longer than DDIM Inversion. TurboEdit (Wu et al., 2024) introduced an encoder-based iterative inversion technique for few-step diffusion models, enabling precise image inversion and disentangled image editing. DirectInv (Ju et al., 2023) separates the source and target diffusion branches for improved text-guided image editing, leading to better content preservation and edit fidelity. AIDI (Pan et al., 2023) and ReNoise (Garibi et al., 2024) used a fixed-point iteration technique at each inversion step to find a better solution for the implicit function posed by DDIM equations and thereby achieving improved accuracy. ExactDPM (Hong et al., 2024), similar to  (Pan et al., 2023; Garibi et al., 2024) proposed gradient-based methods for finding effective solutions for the implicit function.
@@
Alternative methods proposed in (Huberman et al., 2023; Brack et al., 2024; Deutch et al., 2024) use a stochastic DDPM inversion instead of a deterministic one. Although it can recover the exact image during reconstruction, the stochastic nature of these approaches usually causes the method to struggle with reconstructing fine details during editing. These methods require many steps for high-quality editing, making the process time-consuming. Additionally, they demand a larger memory footprint, as they need to store $T+1$ latents for each inverted image, restricting their use only to reconstruction and editing tasks. We also compare our approach with these methods in terms of editing quality and demonstrate that our method outperforms them all.
@@ 3 Preliminaries
$z_{t}=f(z_{t})$
$f(z_{t})$ ::=:= $\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}z_{t-1}+\sqrt{\alpha_{t}}\cdot\Delta\psi(\alpha_{t})\cdot\epsilon_{\theta}(z_{t},t,p).$ (4)
@@
DDIM inversion is fast, but the approximation of Eq. (5) inherently introduces errors at each time step. As these errors accumulate, they cause the whole diffusion process to become inconsistent in the forward and the backward processes, leading to poor image reconstruction and editing (Mokady et al., 2023; Wallace et al., 2023; Pan et al., 2023). This is particularly noticeable in few-step and consistency models with a small number of denoising steps (typically 2-4 steps), where there’s a significant gap between $z_{t}$ and $z_{t-1}$ Garibi et al. (2024).
@@
Iterative inversion and optimization methods. Several papers proposed to improve the approximation using iterative methods. AIDI (Pan et al., 2023) and ReNoise (Garibi et al., 2024) proposed to directly solve Eq. (4) using fixed-point iterations (Burden, 1985), a widely-used method in numerical analysis for solving implicit functions. In a related way, (Hong et al., 2024) solves a more precise inversion equation, obtained by employing higher-order terms, using gradient descent.
Newton-Raphson Root finding. The Newton-Raphson method is a widely used numerical technique for finding roots of a real-valued function $F$ (Burden et al., 2015). It is particularly effective for solving equations of the form $F(z)=0$, when $F:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ for an arbitrary dimension $D$. It provides fast convergence, typically quadratic, by requiring the evaluation of the function and the inversion of its Jacobian matrix. It has also been shown that when initialized near a local extremum, NR may converge to that point (oscillating around it) (Kaw et al., 2003). The NR scheme in general form, is given by
$z^{k+1}_{t}=z^{k}_{t}-J(z^{k}_{t})^{-1}\leavevmode\nobreak\ F(z^{k}_{t}),$ (6)
where $J(z_{t}^{k})^{-1}$ presents the inverse of a Jacobian matrix $J\in\mathbb{R}^{D\times D}$ for $F$ (all derivatives are w.r.t to $z$), and $k$ stands for the iteration number. The iteration starts with an initial guess, $z=z^{0}$.
@@ 4 Our method: Guided Newton Raphson Inversion
Existing inversion methods are either fast-but-inaccurate, like DDIM inversion, or more precise-but-slow, like (Pan et al., 2023) and (Hong et al., 2024). This paper presents a method that achieves a balance of speed and improved precision compared to these existing approaches.
@@ 4.1 Framing inversion as efficient root-finding
and seek the roots $\hat{r}(z_{t})=0$ where |||||$\cdot$||| | ⋅ | | denotes a $L_{1}$ norm which simply sum over all absolute values of $\hat{r}$. Applying the norm on the residual has be done in previous work (Hong et al., 2024; Garibi et al., 2024), however here it is used to reduce the Jacobian to a vector and that can be computed quickly (see derivation in Appendix A). We call this method Newton-Raphson Inversion, NRI.
@@ 4.2 Guided Newton Raphson Inversion
where $\mu_{t}$, $\beta_{t}$ were defined in Section 3, see Eq (1). To guide NR to select an ”in-distribution” solution, we define a new objective $\mathcal{F}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{+}_{0}$:
@@
Here, $\lambda>0$ is a hyperparameter weighting factor for the guidance term. See Appendix E for ablation of $\lambda$ values. Note that this guidance term incorporates prior knowledge about the probable locations of the solutions in $z$ space.
@@
Here $g_{i}:=\frac{\partial\mathcal{F}(z_{t})}{\partial z_{t}(i)}$ is the partial derivative of $\mathcal{F}$ with respect to the variable (descriptor) $z_{t}(i)$. The index $i\in[0,1,...,D]$ indicates $z$’s components and $k$ stands for the iteration number, while $\eta$ is a small constant added for numerical stability. The function $g$ can be computed efficiently using automatic differentiation engines. We initialize the process with $z$ from the previous diffusion time-step $z_{t-1}$. We call our approach GNRI for Guided Newton Raphson Inversion. For discussion on applying NR on our objective function and the guidance term see Appendix B.
@@ 5 Experiments
Compared Methods: We compared our approach with the following methods. Standard DDIM Inversion (Song et al., 2021), SDEdit Meng et al. (2022), Null-text (Mokady et al., 2023) and EDICT (Wallace et al., 2023). Two fixed-point based methods: AIDI (Pan et al., 2023) and ReNoise (Garibi et al., 2024) with further comparison to Anderson Acceleration Pan et al. (2023) shown in Appendix I. Gradient-based method ExactDPM (Hong et al., 2024). Wu2024Turbo (Wu et al., 2024) and two DDPM inversion methods: EditFriendly Huberman et al. (2023) and Gal2024Turbo (Deutch et al., 2024). In all experiments, we used code and parameters provided by the respective authors.
@@ 5.1 Guidance Term Motivation and Comparison to Other Numerical Schemes
To illustrate the motivation behind our guidance term, we compare the reconstruction quality of non-guided NRI a ”vanilla” NR inversion method with two iterative numerical schemes: Fixed-point Iteration (FPI) ((Pan et al., 2023) and Gradient Descent (GD) (Hong et al. (2024). We test the reconstruction of 5,000 image-caption pairs from COCO Lin et al. (2014), presenting three metrics in Fig. 3. Fig 3a shows that FPI, GD and non-guided NRI, all converge to low residual values but the reconstruction values for NRI (green line) in Fig. 3b are notably lower. This suggests that solutions reached by NRI are in areas of the latent space that yield poor results.
@@
Figure 4: (Left) Reconstruction qualitative results: Comparing image inversion-reconstruction performance. While all baseline methods struggle to preserve the original image, GNRI successfully excels in accurately reconstructing it. (Right) Inversion Results: Mean reconstruction quality (y-axis, PSNR) and runtime (x-axis, seconds) on the COCO2017 validation set. Our method achieves high PSNR while reducing inversion-reconstruction time by a factor of ×$\times 2$ (compared to DDIM) and up to ×$\times 40$ (compared to ExactDPM) on SDXL-turbo and ×$\times 10$ to ×$\times 40$ on Flux.1, compared to other approaches.
@@ 5.2 Image Reconstruction
To evaluate the fidelity of our approach, we measure the PSNR of images reconstructed from seed inversions. Specifically, we used the entire set of 5000 images from the MS-COCO-2017 validation dataset (Lin et al., 2014), along with their corresponding captions. For each image-caption pair, we first found the inverted latent and then used it to reconstruct the image using the same caption. Given that the COCO dataset provides multiple captions for each image, we used the first listed caption as a conditioning prompt. Figure 4 (left) demonstrates a qualitative comparison of reconstructed images. Figure 4 (right) illustrates the PSNR of reconstructed images versus inversion time, highlighting the performance of our method compared to SoTA inversion techniques. The dashed black line represents the upper bound set by the Stable Diffusion VAE, showing that our method achieves the highest PSNR among SoTA methods and approaches the VAE upper bound, while also being the fastest. Additional qualitative comparisons for SD2.1 and further analysis are in Appendix H.